OpenStudy (anonymous):

Why \(lim_{n \to \infty}(1-\frac{1}{n})^{n} = e^{-1} \) ? OR How can i find this rule, whats it name ?

OpenStudy (anonymous):

I thought it should be \[\lim{n \to \infty}(1-0)^{n}\]

OpenStudy (anonymous):

hiii, binomial expansion of \[(1+x)^{n}\] substitute for n=1/x you'll get binomial expansion as : \[(1+x)^{1/x}\] = 1+ x*1/x + (1-x)/2! + (1-x)(1-2x)/3!+............ now apply limit x tending to 0 you'll get : 1+1+1/2!+ 1/3!+.......... = e similarly for \[(1-x)^{x}\] we get e-1 :)

OpenStudy (anonymous):

can we find this rule in internet, is it possible to send a link for it ?

OpenStudy (anonymous):

yeah you can google that out !! I answerd that on what I recently studied :)

OpenStudy (anonymous):

ok i think your answer is good enough, but to understand better i need to read once more i think from other source..

OpenStudy (anonymous):

what is this number name \[e^{-1}\] ?

OpenStudy (anonymous):

it's actually exponential to the power -1... we write it as " e " it's value approximately equals to 2.7

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

e-1 = 1/e is it more clear now..?

OpenStudy (anonymous):

ok thank you very much, its clearer

OpenStudy (anonymous):

we can expand e^x as 1+ x/1! + x^2/2! + x^3/3!+......

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

ok thank you meera now its much clearer, thanks for your efforts

OpenStudy (anonymous):

:)